OpenStudy (partyshotz2409):

I will medal What is the length of the hypotenuse of a right triangle if each of the two legs is 7 units? Square root of 14 7 Square root of 98 49

OpenStudy (rayep12):

A

OpenStudy (partyshotz2409):

really

OpenStudy (rayep12):

pretty sure

OpenStudy (partyshotz2409):

no its the sqaure root of 98

OpenStudy (jadeishere):

No, incorrect @rayep12 Pythagoras Theorem; 7^2 + 72 = c^2

OpenStudy (jadeishere):

So 49 + 49 = hypotenuse^2

OpenStudy (partyshotz2409):

i forgot how to do this then remembered

OpenStudy (rayep12):

right

OpenStudy (partyshotz2409):

its 49+49=98

OpenStudy (partyshotz2409):

and 7^2+7^2=_2 its the sqaure root of 98